# Breaking Memory Limits: Gradient Wavelet Transform Enhances LLMs Training

### v1

@@ Abstract
Large language models (LLMs) have shown impressive performance across a range of natural language processing tasks. However, their vast number of parameters introduces significant memory challenges during training, particularly when using memory-intensive optimizers like Adam. Existing memory-efficient algorithms often rely on techniques such as singular value decomposition projection or weight freezing. While these approaches help alleviate memory constraints, they generally produce suboptimal results compared to full-rank updates. In this paper, we investigate the memory-efficient method beyond low-rank training, proposing a novel solution called Gradient Wavelet Transform (GWT), which applies wavelet transforms to gradients in order to significantly reduce the memory requirements for maintaining optimizer states. We demonstrate that GWT can be seamlessly integrated with memory-intensive optimizers, enabling efficient training without sacrificing performance. Through extensive experiments on both pre-training and fine-tuning tasks, we show that GWT achieves state-of-the-art performance compared with advanced memory-efficient optimizers and full-rank approaches in terms of both memory usage and training performance.
keywords:
Large language models, memory-efficient training, wavelet transform.1 Introduction
@@
Large Language Models (LLMs) have made remarkable strides since the release of ChatGPT, a groundbreaking chatbot based on LLM technology [1]. Their promising performance and scalability have led to rapid adoption across a range of interdisciplinary fields, including science [2, 3, 4], medicine [5, 6, 7, 8], and policy-making in biology [9]. The impressive capabilities of LLMs arise from their vast number of parameters and the massive datasets used for training. However, training these models imposes significant demands on the optimizer, with Adam [10] emerging as the most commonly used optimizer due to its fast convergence, robustness, and strong experimental results.
@@
Despite these advantages, Adam’s memory requirements present a major drawback. Specifically, it consumes twice the memory of the model itself (as illustrated in Figure 1), making memory usage a key bottleneck in LLM training. For instance, pre-training a LLaMA 7B [11] model requires approximately 58GB of GPU memory: 14GB for model weights, 14GB for gradients, 2GB for activations, and 28GB for Adam optimizer states under BF16 precision [12]. When the model size exceeds 100B parameters, as with GPT-3 [1] (175B parameters), the additional memory overhead from Adam can exceed 350GB, requiring at least five additional NVIDIA A100 80GB GPUs.
@@
Given the high costs of training LLMs, methods that reduce the batch size or use larger GPUs quickly become unsustainable. To mitigate these issues, many studies have focused on optimizing memory usage during large-scale model training. A promising solution is low-rank training [13, 12], which has made significant progress in alleviating memory constraints. These memory-efficient methods can be broadly classified into two categories: weight-based and gradient-based approaches [14]. Notable examples include Low-Rank Adaptation (LoRA) [13] and Gradient Low-Rank Projection (GaLore) [12].
@@
LoRA leverages the low-rank nature of model parameter updates by freezing the pre-trained model weights and reparameterizing weight updates with trainable low-rank decomposition matrices. Extensive experiments have demonstrated that LoRA is effective in fine-tuning tasks, but its performance in pre-training remains limited. Additionally, the strong low-rank assumptions underlying LoRA often lead to suboptimal outcomes [15, 16]. To address these limitations, Re-LoRA [17] extends LoRA’s applicability to pre-training by periodically updating the frozen model parameters. However, this periodic update process requires full-rank training, which significantly undermines the memory efficiency gains Re-LoRA could otherwise achieve [12].
@@
In contrast to LoRA, which decomposes the update matrix, GaLore reduces the optimizer state memory usage by performing Singular Value Decomposition (SVD) on the gradients, since the memory occupied by optimizer states is directly tied to the gradient size (as shown in Figure 1). GaLore has demonstrated effective performance in both pre-training and fine-tuning tasks. However, its performance still lags behind that of full-rank methods, and it only captures information within the projected subspace, discarding gradient information outside of it. This limitation becomes particularly problematic when using lower ranks or when gradients are noisy, leading to significant convergence issues [12, 18].
@@
To address these challenges, Fira [14] introduces the error residual between the full-rank gradient and its subspace projection, while GoLore [18] enhances GaLore by utilizing random projection to better capture gradient information when noise is dominant. However, these methods still rely on SVD for updating the subspace, and the high computational cost of SVD ( for a matrix of size ) can significantly slow down training, particularly for large models. For example, updating a single subspace may take over ten minutes for the LLaMA 7B model, while inference typically only takes a few seconds [19]. To mitigate this issue, APOLLO [19] proposes an SVD-free variant of GaLore, replacing SVD with random projections. This modification improves both GaLore’s experimental performance and its training speed. Intuitively, the loss of information suggests that low-rank methods should perform worse in training compared to full-rank methods. Surprisingly, Fira and APOLLO outperformed full-rank training in extensive experiments, highlighting the potential of memory-efficient algorithms as alternatives to full-rank methods, which makes memory-efficient training meaningful.
@@
The above methods are based on low-rank decomposition. Aside from low-rank decomposition, Wavelet Transform (WT) [20] also has wide applications in the field of compression. Compared with low-rank decomposition, wavelet transform has lower computational complexity (), better captures local information features, and suppresses high-frequency noise [21]. Moreover, due to its robustness to noise, it is better suited for stochastic gradient descent, a training model method that involves noise. Therefore, we can naturally raise the following question: Can wavelet decomposition, like SVD decomposition, benefit the training of large language models, reducing memory consumption and improving model performance?
In this paper, we explore memory-efficient training methods that extend beyond low-rank techniques. Inspired by the successful use of WT in image compression [22, 23], we introduce Gradient Wavelet Transform (GWT), a plug-and-play framework that applies wavelet transforms to gradients, effectively compressing the dimension and reducing the memory required for storing optimizer states. To illustrate the effectiveness of GWT, we provide a comparison with GaLore and full-rank Adam in terms of memory usage, as shown in Figure 1. Experimentally, we demonstrate that GWT performs well in both pre-training and fine-tuning tasks. Notably, when pre-training LLaMA models on the C4 dataset [24], GWT outperforms full-rank methods, achieving up to a 67% reduction in optimizer memory usage while accelerating training speed with our 2-level GWT on LLaMA 1B. Furthermore, we show that GWT is a versatile framework that can be integrated with optimizers beyond Adam, providing a general solution for memory-efficient training.
2 Results
In this section, we demonstrate that the GWT, as a memory-efficient optimization technique, can achieve performance on par with or even surpassing that of full-rank optimizers, while significantly reducing memory usage and increasing training throughput. Specifically, we evaluate GWT in two contexts: pre-training LLaMA [11] models on the Colossal Clean Crawled Corpus (C4) [24] English benchmark, and fine-tuning the RoBERTa-base model [25] on the General Language Understanding Evaluation (GLUE) [26] benchmark. Additionally, we present an ablation study to explore the impact of hyperparameters, including , learning rate (), and the GWT level .
For the GWT implementation, we employ the discrete Haar and discrete Daubechies-2 (DB2) wavelets as filters. A detailed description of the experimental setup can be found in Appendix A.
2.1 Pre-training LLaMA on C4
We integrate our Gradient Wavelet Transform (GWT) method with the Adam optimizer [10], using default hyperparameters (), to evaluate its performance against existing memory-efficient algorithms on LLaMA models for the C4 pre-training task. For a comprehensive comparison, we include our reproduced results for full-rank Adam [10] and GaLore, and reference results from prior works such as LoRA [13], ReLoRA [17], Fira [14], and APOLLO [19].
| Method |
60M |
130M |
350M |
1B |
| Full-Rank | 33.37 (0.36G) | 25.08 (0.76G) | 18.75 (2.06G) | 16.10 (7.80G) |
Haar-2 |
29.35 |
22.47 |
16.29 |
13.50 |
DB2-2 |
29.60 (0.24G) | 22.59 (0.52G) | 16.30 (1.22G) |
1 |
Haar-3 |
29.81 (0.22G) | 22.63 (0.48G) | 16.35 (1.08G) | - (3.81G) |
DB2-3 |
30.01 (0.22G) | 22.68 (0.48G) | 17.98 (1.08G) | - (3.81G) |
| GaLore-1/8 | 48.48 (0.22G) | 30.02 (0.48G) | 21.59 (1.08G) | - (3.81G) |
| APOLLO∗ [19] | 31.55 (0.24G) | 22.84 (0.52G) | 16.67 (1.22G) | 14.10 (4.38G) |
| Low-Rank∗ [12] | 78.18 (0.26G) | 45.51 (0.54G) | 37.41 (1.08G) | 142.53 (3.57G) |
| ReLoRA∗ [12] | 37.04 (0.36G) | 29.37 (0.80G) | 29.08 (1.76G) | 18.33 (6.17G) |
| Training Tokens | 1.1B | 2.2B | 6.4B | 13.1B |
In all experiments, we use a learning rate of 0.01 for both GWT and GaLore and apply cosine annealing for the learning rate schedule. For GaLore-r, we set the model rank to 1/4 and 1/8 model rank, denoted as GaLore-1/4 and GaLore-1/8, corresponding to the use of a 2/3-level GWT. We follow the initialization strategy from prior works [12] to initialize the network weights and enable GWT in both the Multi-Layer Perceptron (MLP) and attention modules of the model. Haar and DB2 wavelets are used as basis functions, with Haar-2 specifically referring to the optimizer with a 2-level Haar transform.
The final validation perplexity and estimated memory usage for all methods are summarized in Table 1. Experimental results show that GWT consistently achieves lower validation perplexity while reducing memory usage compared to other memory-efficient baselines. In all the pre-training cases, GWT outperforms full-rank optimizers, and GaLore performs relatively closely to GWT. The choice of wavelet filter, whether Haar or DB2, does not have a significant impact on the final experimental results. Specifically, in the LLaMA 1B pre-training experiment, our 2-level GWT method reduces the total memory cost by 43.8% and the optimizer memory cost by 65.7% (from 7.80GB to 4.38GB on LLaMA 1B). Even with the use of 3-level GWT, our method still achieves experimental results that surpass full-rank, while reducing total memory consumption by up to 50%. In the same scenario, the gap between GaLore-1/8 and full-rank is further widened. Details of the memory estimation are presented in Appendix A.3.
In addition to the final evaluation metrics, we present learning curves for the pre-training task on LLaMA 60M-350M with varying memory costs in Figure 2. These curves show that GWT consistently outperforms both full-rank Adam and GaLore across different levels of wavelet transformation. In most cases, GaLore consistently shows a significant gap compared to full-rank Adam, while all GWT methods maintain results surpassing full-rank. Moreover, GWT not only achieves lower final validation perplexity but also accelerates convergence compared to the baselines.
We also extend our method to the Adafactor optimizer [27] and present the corresponding experimental results in Figure 6. These results demonstrate that, compared to GaLore, GWT achieves both lower validation perplexity and performance closer to that of the full-rank optimizer. This highlights the versatility and effectiveness of our method when integrated with other optimizers, showcasing its broad applicability in memory-efficient optimization.
Beyond final validation perplexity and memory usage, we also evaluate token throughput during the pre-training of the LLaMA 1B model using Haar-2, GaLore, and full-rank Adam. The experimental results are summarized in Table 2. Our method (Haar-2) achieves a training throughput of 2.825K tokens per second on GPU, providing more than a 2x speedup on throughput compared to full-rank Adam and approximately a 12% speedup relative to GaLore. This demonstrates that, in addition to its memory efficiency, the GWT method significantly accelerates training throughput, making it a highly efficient alternative to traditional optimization methods for large-scale pre-training tasks.
20K |
40K |
60K |
80K |
Tokens/s |
| Haar-2 | 18.34 | 15.76 | 14.47 | 13.72 | 2.825K |
| GaLore | 20.37 | 17.82 | 16.44 | 15.79 | 2.518K |
| Full-Rank | 21.77 | 18.19 | 16.71 | 16.16 | 1.259K |
| Tokens (B) | 2.6 | 5.2 | 7.8 | 10.4 |
2.2 Fine-Tuning on GLUE Benchmarks
In this section, we further evaluate our method by fine-tuning the RoBERTa-base model on the GLUE benchmark. We report performance metrics across various tasks: accuracy for MNLI, Matthew’s correlation for CoLA, Pearson correlation for STS-B, F1 score for MRPC, and accuracy for the remaining tasks. It is important to note that the optimizer memory ratio during fine-tuning is typically very low, such as 1/96. For GWT, this would require approximately 7 wavelet transformations, which introduces significant computational overhead and contradicts the original intent of GWT. Therefore, for a fair comparison, we use GaLore-1/4 and Haar-2, referencing the results from LoRA and full-rank Adam as reported in prior work [12]. In the fine-tuning task, our primary focus is on evaluating the effectiveness of GWT. There is, however, considerable potential for accelerating multi-level GWT, which we plan to explore in future work.
The experimental results are presented in Table 3. As shown, GWT proves to be effective not only for pre-training but also for fine-tuning, offering competitive performance across multiple GLUE tasks. In contrast, LoRA is primarily designed for fine-tuning tasks and does not exhibit the same versatility across both pre-training and fine-tuning stages as GWT does.
CoLA |
STS-B |
MRPC |
RTE |
SST2 |
MNLI |
QNLI |
QQP |
Avg. |
| Full-Rank∗ | 62.24 | 90.92 | 91.30 |
79.42 |
94.57 |
87.18 | 92.33 |
92.28 |
86.28 |
| Haar-2 |
62.57 |
91.16 |
93.26 |
79.42 |
94.26 |
87.37 |
92.53 | 91.94 |
86.56 |
| GaLore | 61.32 | 91.13 | 92.41 | 77.25 | 94.03 | 86.77 |
92.56 |
91.77 | 85.90 |
2.3 Hyperparameter Study
In this section, we investigate the effects of the hyperparameters and GWT level . A more detailed discussion of these two hyperparameters is provided in Section 4.4.
Scale . We evaluate Adam with 2-level Haar GWT using different values of at a fixed learning rate (). As shown in Figure 3a and 3b, we observe that our GWT is invariant to the hyperparameter. For values larger than , the final experimental results remain relatively stable and do not differ significantly. Based on the network initialization strategy we used, most modules in the network have a learning rate greater than 0.001 (). In contrast to Adam, which often requires a learning rate smaller than 0.001 to avoid loss spikes, GWT shows significantly better robustness to higher learning rates. For fine-tuning tasks, tuning is equivalent to adjusting the learning rate () hyperparameter based on our strategy. A detailed study of the learning rate hyperparameters (,) on fine-tuning tasks is presented in Appendix A.2.
GWT level . We investigate the effects of the GWT level hyperparameter , with the corresponding validation perplexity curves shown in Figure 3c and 3d. We found that the value of has little impact on the final experimental results, and lower values of are only slightly better than others. One question that arises from the experimental results is whether the approximation coefficients are not crucial when using wavelets to compress gradients in LLM training, but rather the detail coefficients. The main impact of seems to be on memory usage and throughput. The throughput results and optimizer memory estimate for different GWT levels are summarized in Table 4, demonstrating the trade-off between memory efficiency and model performance. Higher wavelet levels improve memory usage, but this may come at the cost of slightly higher validation perplexity and slower training speed.
| Methods | 60M | 130M | 350M | |||
| Memory | Tokens/s | Memory | Tokens/s | Memory | Tokens/s | |
| Haar-1 | 0.16G | 95.8K | 0.36G | 46.0K | 0.81G | 12.7K |
| Haar-2 | 0.12G | 91.9K | 0.28G | 45.1K | 0.54G | 12.1K |
| Haar-3 | 0.10G | 90.1K | 0.24G | 44.8K | 0.40G | 11.5K |
| Haar-4 | 0.09G | 84.2K | 0.22G | 43.8K | 0.33G | 10.8K |
| Haar-5 | 0.08G | 83.5K | 0.21G | 41.5K | 0.29G | 9.25K |
3 Discussion
In this paper, we investigate memory-efficient optimization algorithms that extend beyond conventional low-rank decomposition and quantization techniques. Drawing inspiration from the limitations of current optimization methods and the proven success of WT in image and signal processing, we propose a unified framework that integrates GWT into standard optimizers. This framework bridges the gap between gradient compression and wavelet transforms, enabling more efficient training of LLMs by reducing memory usage and accelerating training speed.
To evaluate the effectiveness of our approach, we integrate GWT into the Adam [10] and Adafactor [27] optimizers. Our experimental results show that GWT achieves state-of-the-art performance across multiple stages of the LLM training process and is fully compatible with other optimizers beyond Adam. This gradient compression technique not only improves final validation scores but also enhances throughput, offering a novel memory-performance trade-off that complements existing methods, particularly those based on low-rank optimizations.
@@ 4 Methods
In this section, we first provide an overview of the basic theory behind the wavelet transform. We then explain how the GWT method is applied to the Adam optimizer [10], followed by a discussion of a unified framework that enables the integration of GWT into other optimizers, such as Adafactor [27]. We also cover the associated hyperparameters for GWT.
@@ 4.1 Discrete Haar Wavelet Transform
Wavelet transforms can be categorized into continuous and discrete types. In this paper, we focus on the 1-dimensional Discrete Haar wavelet Transform (DHT) as a case to illustrate our approach. DHT is one of the simplest and most widely used wavelet transforms in signal and image processing due to its low computational cost [34]. The DHT decomposes a signal (or image) into two components: a) Approximation coefficients (Low-frequency): Which represent the smooth or average part of the signal/image. b) Detail coefficients (High-frequency): Capture the difference or finer details of the signal/image.
For example, for an input signal or image represented by the sequence of values , the 1-level DHT (Haar-1) computes the approximation and detail coefficients as follows

### v3

# Wavelet Meets Adam: Compressing Gradients for Memory-Efficient Training

@@ Abstract
Large language models (LLMs) have shown impressive performance across a range of natural language processing tasks. However, their vast number of parameters introduces significant memory challenges during training, particularly when using memory-intensive optimizers like Adam. Existing memory-efficient algorithms often rely on techniques such as singular value decomposition projection or weight freezing. While these approaches help alleviate memory constraints, they generally produce suboptimal results compared to full-rank updates. In this paper, we investigate the memory-efficient method beyond low-rank training, proposing a novel solution called Gradient Wavelet Transform (GWT), which applies wavelet transforms to gradients in order to significantly reduce the memory requirements for maintaining optimizer states. We demonstrate that GWT can be seamlessly integrated with memory-intensive optimizers, enabling efficient training without sacrificing performance. Through extensive experiments on both pre-training and fine-tuning tasks, we show that GWT achieves performance comparable to advanced memory-efficient optimizers and full-rank approaches in terms of both memory usage and training performance.
1 Introduction
@@
Large Language Models (LLMs) have made remarkable strides since the release of ChatGPT [2]. Their promising performance and scalability have led to rapid adoption across a range of interdisciplinary fields, including science [1, 27, 28], medicine [13, 29, 33, 49], and policy-making in biology [20]. The impressive capabilities of LLMs arise from their vast number of parameters and the massive datasets used for training. However, training these models imposes significant demands on the optimizer, with Adam [22] emerging as the most commonly used optimizer due to its fast convergence, robustness, and considerable experimental results.
@@
Despite these advantages, Adam’s memory requirements present a major drawback. Specifically, it consumes twice the memory of the model itself (as illustrated in Figure 1), making memory usage a key bottleneck in LLM training. For instance, pre-training a LLaMA 7B [43] model requires approximately 58GB of GPU memory: 14GB for model weights, 14GB for gradients, 2GB for activations, and 28GB for Adam optimizer states under BF16 precision [54]. When the model size exceeds 100B parameters, as with GPT-3 [2] (175B parameters), the additional memory overhead from Adam can exceed 700GB, requiring at least 9 additional NVIDIA A100 80GB GPUs.
@@
Given the high costs of training LLMs, methods that reduce the batch size or use larger GPUs quickly become unsustainable. To mitigate these issues, many studies have focused on optimizing memory usage in LLM training. A promising solution is low-rank training [21, 54], which has made significant progress in alleviating memory constraints. These memory-efficient methods can be broadly classified into two categories: weight-based and gradient-based approaches [6]. Notable examples include Low-Rank Adaptation (LoRA) [21] and Gradient Low-Rank Projection (GaLore) [54].
@@
LoRA leverages the low-rank nature of model parameter updates by freezing the pre-trained model weights and reparameterizing weight updates with trainable low-rank decomposition matrices. Extensive experiments have demonstrated that LoRA is effective in fine-tuning tasks, but its performance in pre-training remains limited. Additionally, the strong low-rank assumptions underlying LoRA often lead to suboptimal outcomes [50, 52]. FLoRA [16], on the other hand, uses resampling projection matrices to address the low-rank constraints of LoRA.
@@
In contrast to LoRA, which decomposes the update matrix, GaLore reduces the optimizer state memory usage by performing Singular Value Decomposition (SVD) on the gradients, since the memory occupied by optimizer states is directly tied to the gradient size (as shown in Figure 1). GaLore has demonstrated effective performance in both pre-training and fine-tuning tasks. However, its performance still lags behind that of full-rank methods, and it only captures information within the projected subspace, discarding gradient information outside of it. This limitation becomes particularly problematic when using lower ranks or when gradients are noisy [6, 17].
@@
To address these challenges, Fira [6] introduces the error residual between the full-rank gradient and its subspace projection, while GoLore [17] enhances GaLore by utilizing random projection to better capture gradient information when noise is dominant. However, these methods still rely on SVD for updating the subspace, and the high computational cost of SVD ( for a matrix of size ) can significantly slow down training, particularly for large models. To mitigate this issue, APOLLO [56] proposes an SVD-free variant of GaLore, replacing SVD with random projections. This modification not only improves training speed but also leads to enhanced empirical performance.
While it is intuitive to expect that information loss in low-rank approximations would degrade performance relative to full-rank training, extensive experiments have shown otherwise: both Fira and APOLLO have surprisingly outperformed full-rank training in various settings. These findings underscore the potential of memory-efficient optimization algorithms as strong alternatives to full-rank methods, making efficient training not only necessary but also advantageous in practice.
@@
The above methods are based on low-rank decomposition. Aside from low-rank decomposition, the Wavelet Transform (WT) also has wide applications in the field of compression. Compared with low-rank decomposition, wavelet transform has lower computational complexity (), better captures local information features, and suppresses high-frequency noise [37]. Moreover, due to its robustness to noise, it is better suited for stochastic gradient descent, a training model method that involves noise. Therefore, we can naturally raise the following question:
To address the limitations of existing low-rank training methods, we introduce Gradient Wavelet Transform (GWT), a simple yet effective plug-and-play framework that applies wavelet transforms to gradients during training. By leveraging the natural sparsity and localization properties of wavelets, GWT compresses gradient representations and significantly reduces the memory required to store optimizer states without relying on low-rank assumptions. Distinct from traditional matrix-based compression techniques, GWT is the first method to adopt a frequency-domain perspective for gradient compression. It achieves memory efficiency by selectively controlling low- and high-frequency coefficients. When applied to large-scale language model pretraining, such as LLaMA on the C4 dataset, GWT achieves up to 71% optimizer memory reduction and 1.9× training speedup on 3B models, all while maintaining or even improving model quality. Moreover, GWT integrates seamlessly with optimizers beyond Adam, highlighting its versatility across different training regimes. These results demonstrate that GWT provides a practical, optimizer-agnostic, and highly efficient solution for memory-constrained deep learning.
2 Related Works
Memory-Efficient Optimizers. Given the high memory requirements of the Adam optimizer [22], numerous approaches aim to improve the optimizer or create new optimization methods to reduce training costs for LLMs. Low-Rank Adaptation (LoRA) [21], which addresses this issue by freezing the parameters of the pre-trained model and introducing trainable low-rank matrices that decompose the pre-trained weights. LoRA has proven to be an effective strategy for optimizing the training of large language models, and several subsequent studies have built upon this framework to further enhance optimization performance [16, 23, 50, 52]. Gradient Low-Rank Projection (GaLore) [54] lowers the memory cost of optimizer states by projecting gradients into a lower-dimensional subspace. This method not only facilitates full-parameter LLM training but also delivers strong performance in both pretraining and fine-tuning tasks. Recent GaLore-based extensions have shown experimental results that outperform full-rank methods [6, 56].
System Memory Efficiency. Beyond optimizing memory usage at the level of the optimizer, some strategies target reducing memory consumption in the overall system. Gradient checkpointing [5] is one such method that conserves memory by not storing activation values. Memory offloading [35] improves memory availability during training by utilizing non-GPU memory resources. Quantization techniques [8, 9] further reduce memory overhead by employing lower-bit data representations. These methods have significantly accelerated and optimized LLM training processes. Our GWT approach offers a distinct advantage over these techniques by providing a memory-efficient alternative. Moreover, our method is designed to be compatible with existing system-level optimization strategies, allowing for a synergistic integration that can lead to even faster and more efficient LLM training.
@@ 3 Methods
In this section, we first provide an overview of the fundamental theory behind wavelet transforms, followed by a detailed explanation of how the proposed Gradient Wavelet Transform (GWT) method is integrated into the Adam optimizer [22].
@@ 3.1 Discrete Haar Wavelet Transform
Wavelet transforms can be categorized into continuous and discrete types. In this paper, we mainly focus on the discrete Haar wavelet due to its simple and easy-to-understand basis functions, high computational efficiency, broad applicability to compression tasks [32], and the fact that the total amount of data after decomposition equals the original data, without introducing extra information. It decomposes a signal (or image) into two main components: (a) Approximation coefficients (low-frequency), which represent the smooth or average part of the signal/image, and (b) Detail coefficients (high-frequency), which capture the finer details or differences in the signal/image.
For example, for an input signal or image represented by the sequence of values , the 1-level Discrete Haar Transform (DHT, Haar-1) computes the Approximation and Detail coefficients as follows